Main INDEX
Monthly INDEX
PREV
NEXT
Make New Entry,
Make Followup Entry
User name V. Sulkosky
Log entry time 18:54:45 on February 25, 2009
Entry number 263312
This entry is a followup to: 263006
Followups:
keyword=Experiment Summary for Feb 23, Swing - Feb 24, Swing
Summary for Feb 23 swing - Feb 24 swing
Smooth data taking at fourth pass.
Beam in Use
shift |
ABU |
BANU |
BNA |
ACC |
2009/02/23 swing | 5.49 |
1.11 | 0.69 | 0.7 |
2009/02/24 owl | 7.55 |
0.03 | 0.43 | 0 |
2009/02/24 day | 5.59 |
0.92 | 1.46 | 0 |
2009/02/24 swing | 6.38 |
0.33 | 1.3 | 0 |
total | 25.0 |
2.4 | 3.9 | 0.7 |
Main events/issues:
- February 23rd, Swing shift
- Unpolarized cross section measurements for left HRS.
- Target spin flip enabled every 20 minutes.
- 16:00-18:30 Beam checkout procedure after pass change.
- 18:30 Begin kinematic 4K: Plhrs = 1.6 GeV/c,
45o.
- February 24th
- Production for kinematic 4K.
- 00:25 VESDA AHU alarm, blinking red at the alert level.
- 09:00 Controlled access to investigate cause for VESDA alert.
- 15:45-16:45 MCC RF problem.
- 22:03 Completed kinematic 4K,
change to kinematic 4J: Plhrs = 1.51 GeV/c,
45o.
- 22:45 Begin production on 3He for kinematic 4J.
Plans for the next couple of days:
- Production for kinematic 4J and 4C until 9 am on February
26th.
- February 26th from 09:00 to 15:00
- Beam studies
- Controlled access for spectrometer magnet work
- Pass change to 5th, wien angle to 30o
- Resume polarized production at 5th pass.
A copy of this log entry has been emailed to: brads,jpchen,jiang,meziani,choi